# PEL-NAS: Search Space Partitioned Architecture Prompt Co-Evolutionary LLM-driven Hardware-Aware Neural Architecture Search

### v2

@@ Abstract
Hardware-Aware Neural Architecture Search (HW-NAS) requires joint optimization of accuracy and latency under device constraints. Traditional supernet-based methods require multiple GPU days per dataset. Large Language Model (LLM)-driven approaches avoid training a large supernet and can provide quick feedback, but we observe an exploration bias: the LLM repeatedly proposes neural network designs within limited search space and fails to discover architectures across different latency ranges in the entire search space. To address this issue, we propose PEL-NAS: a search space Partitioned, architecture prompt co-Evolutionary and LLM-driven Neural Architecture Search that can generate neural networks with high accuracy and low latency with reduced search cost. Our proposed PEL-NAS has three key components: 1) a complexity-driven partitioning engine that divides the search space by complexity to enforce diversity and mitigate exploration bias; 2) an LLM-powered architecture prompt co-evolution operator, in which the LLM first updates a knowledge base of design heuristics based on results from the previous round, then performs a guided evolution algorithm on architectures with prompts that incorporate this knowledge base. Prompts and designs improve together across rounds which avoids random guesswork and improve efficiency; 3) a zero-cost predictor to avoid training a large number of candidates from scratch. Experimental results show that on HW-NAS-Bench, PEL-NAS can achieve overall higher HV, lower IGD, and up to 54% lower latency than baselines at similar accuracy. Meanwhile, the search cost drops from days to minutes compared with traditional supernet baselines.
@@ 1 Introduction
Figure 1: Comparison of three generation strategies on HW-NAS-Bench (Edge GPU, CIFAR-10): normal prompt (orange), latency-optimized prompt (blue), and PEL-NAS (green). Latency-optimized prompting increases coverage compared to standard prompting but still leaves gaps, while PEL-NAS achieves near-complete coverage across latency ranges
@@
To address the above two challenges, we propose PEL-NAS: a search space Partitioned, architecture prompt co-Evolutionary and LLM-driven Neural Architecture Search (Figure 2), to reduce exploration bias while improving search efficiency. Our approach begins with a a complexity-driven partitioning strategy that decomposes the vast search space into subspaces with different complexity or different parameter size levels. With the partitioning strategy, PEL-NAS can discover subnetworks across the whole search space, as shown in Figure 1(c). Within each subspace, we then employ an LLM-Powered Evolutionary Operator that functions as an expert reasoning engine, guided by a continually refined Co-evolve Knowledge Base. For each new design, the LLM provides a detailed rationale for its modifications, and a rapid, training-free evaluation protocol provides instant feedback. This synergistic framework transforms the search from a biased, unconstrained generation task into a structured, diverse, and efficient exploration. With our method, we obtain a more complete and dominant Pareto front of hardware-optimized models, achieving near-perfect quality scores. This is accomplished while dramatically reducing the search cost from multiple GPU-days, typical for supernet-based approaches, to mere minutes. The contributions are summarized as follows:
@@
  • Compared to conventional and unconstrained LLM-driven methods, our training-free framework discovers a more complete and dominant set of optimal trade-offs. This superiority is validated by two standard metrics: a significantly higher Hypervolume (HV), indicating our solutions achieve broader coverage of the performance space with both superior and more diverse models, and a lower Inverted Generational Distance (IGD), showing our discovered architectures are closer to the true optimal front. The experiments demonstrate that PEL-NAS enables this with a search cost of minutes, in stark contrast to the days of GPU training required by supernet-based approaches.
@@ 3 Methodology
Figure 2: The PEL-NAS framework. The search space is partitioned into complexity-based niches, where an LLM performs parallel evolutionary search. The individual results are then aggregated to form the final, complete Pareto front, mitigating exploration bias
Our method, PEL-NAS, overcomes the critical exploration bias of LLMs in HW-NAS while preserving the efficiency of training-free methods. As illustrated in Figure 2, our approach integrates three key components: a search space partitioning strategy to ensure diversity, an LLM-powered evolutionary engine for intelligent exploration, and a training-free evaluator to provide rapid feedback on accuracy and latency.
@@ 3.1 Complexity-Driven Search Space Partitioning
Figure 3: Analysis of the HW-NAS-Bench search space. The distribution of total parameters exhibits clear clustering, where each cluster corresponds to a specific number of nor_conv_3x3
@@ 3.2 LLM-Powered Partitioned Co-Evolution of Prompts and Architectures
Figure 4: The Co-evolve Prompt Generator in PEL-NAS. The LLM first acts as a reasoning engine, updating a Knowledge Base by analyzing prior results. This learned knowledge then informs the LLM’s second role as an expert architect, where it generates new, rationale-driven architectures under specific constraints, creating a self-optimizing search process
@@
Knowledge Base Update After each search cycle, PEL-NAS collects the architectures along with their measured accuracy, latency and the corresponding design rationales from previous cycle. The LLM first acts as a reasoning engine, analyzing these results to update a Co-evolve Knowledge Base. For example, the Knowledge Base may update rules such as “avg_pool takes a long time and has limited accuracy improvement” and delete “avg_pool always improves accuracy”. By continuously summarizing such patterns, the LLM accumulates long-term memory of effective design principles and avoids repeatedly exploring unpromising regions, preventing local mode collapse.
@@ 4.1 Experimental Setup
Baselines. We position PEL-NAS against a diverse set of state-of-the-art NAS methods to highlight its unique advantages. Our comparison includes influential supernet-based methods that do not primarily focus on hardware constraints, such as the classic differentiable approach DARTS Liu et al. (2018) and the fairness-enforcing FairNAS Chu et al. (2021). To benchmark against a hardware-aware contemporary, we include PRP-NAS Benmeziane et al. (2023), which represents supernet methods that explicitly optimize for hardware efficiency. Furthermore, we contrast our approach with the latest advancements in LLM-driven search by LLMatic Nasir et al. (2024), that also utilize large language models for architecture generation but are not designed with hardware awareness as a primary objective. For ViT on ImageNet-1k, we report ViT-B/16 Dosovitskiy et al. (2020), DeiT-B Touvron et al. (2021), and the NAS search method AutoFormer Chen et al. (2021).
@@ 4.2 Main Results
Our main experimental results are detailed in Tables 2,  3 and  4. Collectively, they demonstrate that PEL-NAS not only discovers architectures that achieve a balance between accuracy and latency, a Pareto front of superior quality and completeness, but also achieves this with unparalleled efficiency.
Table 3: HV and IGD comparison on HW-NAS-Bench across six hardware devices on CIFAR-10, CIFAR-100, and ImageNet16-120. PEL-NAS consistently outperforms all baselines, demonstrating its ability to find a more complete and dominant Pareto front. (Higher HV is better, lower IGD is better). Best results are in bold
@@
PEL-NAS 0.953 0.043 0.988 0.011 0.943 0.033 0.983 0.042 0.945 0.050 0.972 0.028
Analysis of Discovered Architectures of HW-NAS-Bench on CIFAR-10. Beyond the overall front quality, the individual architectures in Table 2 highlight the practical value of our method. PEL-NAS not only finds models with state-of-the-art accuracy, matching the performance of costly supernet-based methods, but also excels in the low-latency domain where other approaches falter. Crucially, it discovers the undisputed fastest architecture for each hardware target. For example, it identifies a model with a latency of just 1.78ms on the Edge GPU and 1.65ms on the FPGA which outperforming PRP-NAS-BL by over 22% and 54% respectively. This proves its superior ability to explore the full spectrum of trade-offs and deliver a truly comprehensive set of optimal solutions.
Pareto Front Quality Evaluation with HV and IGD. The core quantitative results in Table 3 compare the discovered Pareto fronts using HV and IGD. Across all three datasets and six hardware targets, PEL-NAS consistently and significantly outperforms all baselines. PEL-NAS achieves higher HV scores and lower IGD scores compared to baselines. For example, on CIFAR-10, PEL-NAS can achieve up to 80.6% higher HV and 53.6% lower IGD compared to the non-constrained LLM method. On CIFAR-100, PEL-NAS outperforms LLMatic, FairNAS, PRP-NAS, by 46.5%, 5.7%, 17.4% in HV, and by 34.9%, 6.1%, 18.6% in IGD, in average respectively. These observations further confirm that the front discovered by PEL-NAS is not only larger in volume but also much closer to the true optimal front. The experimental results also demonstrate that our complexity-driven partitioning strategy is highly effective in mitigating the LLM’s generative bias and enabling a more complete and diverse exploration of the search space.
@@
PEL-NAS (Ours) 3 mins (API Calls)
Search Cost. Crucially, as shown in Table 4, PEL-NAS achieves these results with negligible computational cost. As a training-free method, its search cost is measured in API calls (120 times) and minutes, starkly contrasting with supernet-based methods like FairNAS Chu et al. (2021) that require days of GPU training. In contrast, LLMatic Nasir et al. (2024) is the most time-consuming because it needs to train every generated architecture from scratch. This combination of superior search capability and extreme efficiency makes PEL-NAS a practical and powerful solution for real-world HW-NAS challenges.
@@ 4.3 Ablation Studies
Table 5: Ablation study results on CIFAR-100 showing the impact of each component of PEL-NAS. Both the partitioning strategy and the ZC ensemble are shown to be critical components, with their removal causing the most significant performance degradation
@@
To isolate the contribution of each key component of our framework, we conduct a series of ablation studies. The aggregated results are summarized in Table 5, while detailed line graphs illustrating the search process for three datasets across six devices are available in the Appendix (Figures 6, 7, and 8). The analysis reveals that the partitioning strategy is the most critical element. Removing it (- without Partitioning) leads to a catastrophic performance collapse, which provides direct evidence that our niching approach is essential for mitigating the LLM’s mode collapse. Similarly, the ZC ensemble predictor is vital; replacing it with a single Synflow proxy (- without ZC Ensemble) causes a significant performance degradation, confirming that a robust performance signal is crucial to guide the search effectively. Finally, while the partitioned evolutionary algorithm (PEA) (- without LLM Operator) still performs well, it is clearly surpassed by the full PEL-NAS model. This demonstrates that the LLM acts as an intelligent operator, leveraging context to generate superior candidates and further enhancing search efficiency.
@@ 4.4 Generalizability on Vision Transformer Search Spaces
To validate the generalizability of PEL-NAS beyond CNNs, we extend our framework to a Vision Transformer (ViT) search space derived from AutoFormer Chen et al. (2021). We conduct our hardware-aware search experiments on the ImageNet dataset. To ensure an efficient search process, we employ an accuracy predictor. Specifically, we adopt the Auto-Proxy predictor from ViT-Bench-101 Wei et al. (2024), which achieves a strong Spearman’s rank correlation of $91.01\pm 2.63$} on this task, confirming its reliability for performance estimation. All reported accuracies and the resulting Pareto front in Figure 5 are based on the outputs of this predictor.
Figure 5: The Pareto front discovered by PEL-NAS for three AutoFormer search spaces on ImageNet. Latency is evaluated using a single NVIDIA A6000 GPU, and accuracy is estimated via a predictor
To create a realistic hardware-aware scenario, we profile the latency of each candidate architecture directly on our target device, a single NVIDIA A6000 GPU. We then apply the core principle of PEL-NAS—complexity-driven partitioning. Our analysis of the ViT architecture (see Appendix C for a detailed breakdown) reveals that computational complexity, a strong proxy for latency, is dominated by two key parameters: Embed Dim (quadratic impact, $O(D^{2})$) and Depth Num (linear impact, $O(L)$). These parameters govern the scale of the MLP and the number of blocks, respectively, making them the most influential factors. We therefore partition the search space into niches based on discrete ranges of Embedding Dimension and Depth Number, enabling the LLM to efficiently explore trade-offs within structurally similar architectural families. The results, depicted in Figure 5 and detailed in Table 6, underscore the efficacy of our approach. PEL-NAS successfully identifies a dominant Pareto front, discovering architectures with superior accuracy-latency trade-offs.
Table 6: Comparison of Vision Transformer models found by PEL-NAS against state-of-the-art NAS methods. Latency is measured on A6000 GPU
@@
PEL-NAS-ViT-Tiny (Ours) 76.2 4.0 6.9
PEL-NAS-ViT-Small (Ours) 79.7 4.7 16.1
PEL-NAS-ViT-Base (Ours) 82.5 5.4 20.2
@@ 5 Conclusion
In this work, we introduce PEL-NAS, a novel training-free framework designed to counteract the exploration bias inherent in LLM-driven neural architecture search. Our core contribution is a complexity-driven partitioning strategy that divides the search space into distinct niches, compelling the LLM to act as a parallel evolutionary engine and structurally enforcing population diversity across the entire architectural complexity spectrum. This approach effectively mitigates the LLM’s tendency to converge on a narrow set of familiar architectures. Extensive experiments on HW-NAS-Bench demonstrate that PEL-NAS discovers a more complete and dominant Pareto front than baseline methods, validated by significantly superior HV and IGD scores. Our findings present a new paradigm for harnessing LLMs in combinatorial optimization, suggesting that imposing structural constraints on the generative process is a powerful method for mitigating inherent biases, future work could focus on automating the partitioning strategy and applying this framework to other complex design domains.

### v3

# LLM-NAS: LLM-driven Hardware-Aware Neural Architecture Search

@@ Abstract
Hardware-Aware Neural Architecture Search (HW-NAS) requires joint optimization of accuracy and latency under device constraints. Traditional supernet-based methods require multiple GPU days per dataset. Large Language Model (LLM)-driven approaches avoid training a large supernet and can provide quick feedback, but we observe an exploration bias: the LLM repeatedly proposes neural network designs within limited search space and fails to discover architectures across different latency ranges in the entire search space. To address this issue, we propose LLM-NAS: an LLM-driven Neural Architecture Search that can generate neural networks with high accuracy and low latency with reduced search cost. Our proposed LLM-NAS has three key components: 1) a complexity-driven partitioning engine that divides the search space by complexity to enforce diversity and mitigate exploration bias; 2) an LLM-powered architecture prompt co-evolution operator, in which the LLM first updates a knowledge base of design heuristics based on results from the previous round, then performs a guided evolution algorithm on architectures with prompts that incorporate this knowledge base. Prompts and designs improve together across rounds which avoids random guesswork and improve efficiency; 3) a zero-cost predictor to avoid training a large number of candidates from scratch. Experimental results show that on HW-NAS-Bench, LLM-NAS can achieve overall higher HV, lower IGD, and up to 54% lower latency than baselines at similar accuracy. Meanwhile, the search cost drops from days to minutes compared with traditional supernet baselines.
@@ 1 Introduction
Figure 1: Comparison of three generation strategies on HW-NAS-Bench (Edge GPU, CIFAR-10): normal prompt (orange), latency-optimized prompt (blue), and LLM-NAS (green). Latency-optimized prompting increases coverage compared to standard prompting but still leaves gaps, while LLM-NAS achieves near-complete coverage across latency ranges
@@
To address the above two challenges, we propose LLM-NAS: a search space Partitioned, architecture prompt co-Evolutionary and LLM-driven Neural Architecture Search (Figure 2), to reduce exploration bias while improving search efficiency. Our approach begins with a a complexity-driven partitioning strategy that decomposes the vast search space into subspaces with different complexity or different parameter size levels. With the partitioning strategy, LLM-NAS can discover subnetworks across the whole search space, as shown in Figure 1(c). Within each subspace, we then employ an LLM-Powered Evolutionary Operator that functions as an expert reasoning engine, guided by a continually refined Co-evolve Knowledge Base. For each new design, the LLM provides a detailed rationale for its modifications, and a rapid, training-free evaluation protocol provides instant feedback. This synergistic framework transforms the search from a biased, unconstrained generation task into a structured, diverse, and efficient exploration. With our method, we obtain a more complete and dominant Pareto front of hardware-optimized models, achieving near-perfect quality scores. This is accomplished while dramatically reducing the search cost from multiple GPU-days, typical for supernet-based approaches, to mere minutes. The contributions are summarized as follows:
@@
  • Compared to conventional and unconstrained LLM-driven methods, our training-free framework discovers a more complete and dominant set of optimal trade-offs. This superiority is validated by two standard metrics: a significantly higher Hypervolume (HV), indicating our solutions achieve broader coverage of the performance space with both superior and more diverse models, and a lower Inverted Generational Distance (IGD), showing our discovered architectures are closer to the true optimal front. The experiments demonstrate that LLM-NAS enables this with a search cost of minutes, in stark contrast to the days of GPU training required by supernet-based approaches.
@@ 3 Methodology
Figure 2: The LLM-NAS framework. The search space is partitioned into complexity-based niches, where an LLM performs parallel evolutionary search. The individual results are then aggregated to form the final, complete Pareto front, mitigating exploration bias
Our method, LLM-NAS, overcomes the critical exploration bias of LLMs in HW-NAS while preserving the efficiency of training-free methods. As illustrated in Figure 2, our approach integrates three key components: a search space partitioning strategy to ensure diversity, an LLM-powered evolutionary engine for intelligent exploration, and a training-free evaluator to provide rapid feedback on accuracy and latency.
@@ 3.1 Complexity-Driven Search Space Partitioning
Figure 3: Analysis of the HW-NAS-Bench search space. The distribution of total parameters exhibits clear clustering
@@ 3.2 LLM-Powered Partitioned Co-Evolution of Prompts and Architectures
Figure 4: The Co-evolve Prompt Generator in LLM-NAS. The LLM first acts as a reasoning engine, updating a Knowledge Base by analyzing prior results. This learned knowledge then informs the LLM’s second role as an expert architect, where it generates new, rationale-driven architectures under specific constraints, creating a self-optimizing search process
@@
Knowledge Base Update After each search cycle, LLM-NAS collects the architectures along with their measured accuracy, latency and the corresponding design rationales from previous cycle. The LLM first acts as a reasoning engine, analyzing these results to update a Co-evolve Knowledge Base. For example, the Knowledge Base may update rules such as “avg_pool takes a long time and has limited accuracy improvement” and delete “avg_pool always improves accuracy”. By continuously summarizing such patterns, the LLM accumulates long-term memory of effective design principles and avoids repeatedly exploring unpromising regions, preventing local mode collapse.
@@ 4.1 Experimental Setup
Baselines. We position LLM-NAS against a diverse set of state-of-the-art NAS methods to highlight its unique advantages. Our comparison includes influential supernet-based methods that do not primarily focus on hardware constraints, such as the classic differentiable approach DARTS Liu et al. (2018) and the fairness-enforcing FairNAS Chu et al. (2021). To benchmark against a hardware-aware contemporary, we include PRP-NAS Benmeziane et al. (2023), which represents supernet methods that explicitly optimize for hardware efficiency. Furthermore, we contrast our approach with the latest advancements in LLM-driven search by LLMatic Nasir et al. (2024), that also utilize large language models for architecture generation but are not designed with hardware awareness as a primary objective. For ViT on ImageNet-1k, we report ViT-B/16 Dosovitskiy et al. (2020), DeiT-B Touvron et al. (2021), and the NAS search method AutoFormer Chen et al. (2021).
@@ 4.2 Main Results
Our main experimental results are detailed in Tables 2,  3 and  4. Collectively, they demonstrate that LLM-NAS not only discovers architectures that achieve a balance between accuracy and latency, a Pareto front of superior quality and completeness, but also achieves this with unparalleled efficiency.
Table 3: HV and IGD comparison on HW-NAS-Bench across six hardware devices on CIFAR-10, CIFAR-100, and ImageNet16-120. LLM-NAS consistently outperforms all baselines, demonstrating its ability to find a more complete and dominant Pareto front. (Higher HV is better, lower IGD is better). Best results are in bold
@@
LLM-NAS 0.953 0.043 0.988 0.011 0.943 0.033 0.983 0.042 0.945 0.050 0.972 0.028
Analysis of Discovered Architectures of HW-NAS-Bench on CIFAR-10. Beyond the overall front quality, the individual architectures in Table 2 highlight the practical value of our method. LLM-NAS not only finds models with state-of-the-art accuracy, matching the performance of costly supernet-based methods, but also excels in the low-latency domain where other approaches falter. Crucially, it discovers the undisputed fastest architecture for each hardware target. For example, it identifies a model with a latency of just 1.78ms on the Edge GPU and 1.65ms on the FPGA which outperforming PRP-NAS-BL by over 22% and 54% respectively. This proves its superior ability to explore the full spectrum of trade-offs and deliver a truly comprehensive set of optimal solutions.
Pareto Front Quality Evaluation with HV and IGD. The core quantitative results in Table 3 compare the discovered Pareto fronts using HV and IGD. Across all three datasets and six hardware targets, LLM-NAS consistently and significantly outperforms all baselines. LLM-NAS achieves higher HV scores and lower IGD scores compared to baselines. For example, on CIFAR-10, LLM-NAS can achieve up to 80.6% higher HV and 53.6% lower IGD compared to the non-constrained LLM method. On CIFAR-100, LLM-NAS outperforms LLMatic, FairNAS, PRP-NAS, by 46.5%, 5.7%, 17.4% in HV, and by 34.9%, 6.1%, 18.6% in IGD, in average respectively. These observations further confirm that the front discovered by LLM-NAS is not only larger in volume but also much closer to the true optimal front. The experimental results also demonstrate that our complexity-driven partitioning strategy is highly effective in mitigating the LLM’s generative bias and enabling a more complete and diverse exploration of the search space.
@@
LLM-NAS (Ours) 3 mins (API Calls)
Search Cost. Crucially, as shown in Table 4, LLM-NAS achieves these results with negligible computational cost. As a training-free method, its search cost is measured in API calls (120 times) and minutes, starkly contrasting with supernet-based methods like FairNAS Chu et al. (2021) that require days of GPU training. In contrast, LLMatic Nasir et al. (2024) is the most time-consuming because it needs to train every generated architecture from scratch. This combination of superior search capability and extreme efficiency makes LLM-NAS a practical and powerful solution for real-world HW-NAS challenges.
@@ 4.3 Ablation Studies
Table 5: Ablation study results on CIFAR-100 showing the impact of each component of LLM-NAS. Both the partitioning strategy and the ZC ensemble are shown to be critical components, with their removal causing the most significant performance degradation
@@
To isolate the contribution of each key component of our framework, we conduct a series of ablation studies. The aggregated results are summarized in Table 5, while detailed line graphs illustrating the search process for three datasets across six devices are available in the Appendix (Figures 6, 7, and 8). The analysis reveals that the partitioning strategy is the most critical element. Removing it (- without Partitioning) leads to a catastrophic performance collapse, which provides direct evidence that our niching approach is essential for mitigating the LLM’s mode collapse. Similarly, the ZC ensemble predictor is vital; replacing it with a single Synflow proxy (- without ZC Ensemble) causes a significant performance degradation, confirming that a robust performance signal is crucial to guide the search effectively. Finally, while the partitioned evolutionary algorithm (PEA) (- without LLM Operator) still performs well, it is clearly surpassed by the full LLM-NAS model. This demonstrates that the LLM acts as an intelligent operator, leveraging context to generate superior candidates and further enhancing search efficiency.
@@ 4.4 Generalizability on Vision Transformer Search Spaces
To validate the generalizability of LLM-NAS beyond CNNs, we extend our framework to a Vision Transformer (ViT) search space derived from AutoFormer Chen et al. (2021). We conduct our hardware-aware search experiments on the ImageNet dataset. To ensure an efficient search process, we employ an accuracy predictor. Specifically, we adopt the Auto-Proxy predictor from ViT-Bench-101 Wei et al. (2024), which achieves a strong Spearman’s rank correlation of $91.01\pm 2.63$} on this task, confirming its reliability for performance estimation. All reported accuracies and the resulting Pareto front in Figure 5 are based on the outputs of this predictor.
Figure 5: The Pareto front discovered by LLM-NAS for three AutoFormer search spaces on ImageNet. Latency is evaluated using a single NVIDIA A6000 GPU, and accuracy is estimated via a predictor
To create a realistic hardware-aware scenario, we profile the latency of each candidate architecture directly on our target device, a single NVIDIA A6000 GPU. We then apply the core principle of LLM-NAS—complexity-driven partitioning. Our analysis of the ViT architecture (see Appendix C for a detailed breakdown) reveals that computational complexity, a strong proxy for latency, is dominated by two key parameters: Embed Dim (quadratic impact, $O(D^{2})$) and Depth Num (linear impact, $O(L)$). These parameters govern the scale of the MLP and the number of blocks, respectively, making them the most influential factors. We therefore partition the search space into niches based on discrete ranges of Embedding Dimension and Depth Number, enabling the LLM to efficiently explore trade-offs within structurally similar architectural families. The results, depicted in Figure 5 and detailed in Table 6, underscore the efficacy of our approach. LLM-NAS successfully identifies a dominant Pareto front, discovering architectures with superior accuracy-latency trade-offs.
Table 6: Comparison of Vision Transformer models found by LLM-NAS against state-of-the-art NAS methods. Latency is measured on A6000 GPU
@@
LLM-NAS-ViT-Tiny (Ours) 76.2 4.0 6.9
LLM-NAS-ViT-Small (Ours) 79.7 4.7 16.1
LLM-NAS-ViT-Base (Ours) 82.5 5.4 20.2
@@ 5 Conclusion
In this work, we introduce LLM-NAS, a novel training-free framework designed to counteract the exploration bias inherent in LLM-driven neural architecture search. Our core contribution is a complexity-driven partitioning strategy that divides the search space into distinct niches, compelling the LLM to act as a parallel evolutionary engine and structurally enforcing population diversity across the entire architectural complexity spectrum. This approach effectively mitigates the LLM’s tendency to converge on a narrow set of familiar architectures. Extensive experiments on HW-NAS-Bench demonstrate that LLM-NAS discovers a more complete and dominant Pareto front than baseline methods, validated by significantly superior HV and IGD scores. Our findings present a new paradigm for harnessing LLMs in combinatorial optimization, suggesting that imposing structural constraints on the generative process is a powerful method for mitigating inherent biases, future work could focus on automating the partitioning strategy and applying this framework to other complex design domains.